OpenStudy (jkbo):

Please help its a chem question An infrared wave is measured to have a wavelength of 5.6 x 10^4 nm. What is its frequency?

OpenStudy (anonymous):

for electrromagnetic radiation (in a vacuum) \[ c = f\lambda \\ \ \\ c = \ \textrm{speed of light} = 3.0 x 10^8 m/s \\ f = \ \textrm{frequency in Hertz (1/s)} \\ \lambda = \ \textrm{wavelength (in meters)} \]

OpenStudy (jkbo):

@AllTehMaffs alright it would be f = frequency I need help solving it

OpenStudy (anonymous):

\[ f = \frac{c}{\lambda} \] what's your wavelength in meters?

OpenStudy (jkbo):

5.4 x10^12

OpenStudy (jkbo):

@AllTehMaffs

OpenStudy (anonymous):

... no. should be very small

OpenStudy (jkbo):

I'm not sure than @AllTehMaffs

OpenStudy (anonymous):

5.6x10^4 nm 56000nm ( 1x 10^-9)m/nm = 5.6x10^-5 m

OpenStudy (jkbo):

oo ok thank you could you help me with one more? @AllTehMaffs

OpenStudy (anonymous):

sure ^^ Did you finish that one? c / that number?

OpenStudy (jkbo):

Almost all commercially available microwave ovens employ radiation with a frequency of 2.45 x 109 Hz. Calculate the wavelength of this radiation.

OpenStudy (jkbo):

@AllTehMaffs yes I got 1.7857143e-13

OpenStudy (anonymous):

that's really, really small. It should be pretty large. did you divide that number by c instead? it should be about 10^12

OpenStudy (anonymous):

or it should be 10^13

OpenStudy (anonymous):

not -13

OpenStudy (jkbo):

yes I did c and thats what I got

OpenStudy (jkbo):

@AllTehMaffs

OpenStudy (anonymous):

(3 x 10^8 / 5.6 x 10^-5) = 5.36 x 10^12 1/s I'm so sorry, that's what you got originally. I couldn't tell what thing you were talking about without units

OpenStudy (jkbo):

its fine haha @AllTehMaffs

OpenStudy (anonymous):

^^ what's your next question?

OpenStudy (jkbo):

Almost all commercially available microwave ovens employ radiation with a frequency of 2.45 x 109 Hz. Calculate the wavelength of this radiation. @AllTehMaffs

OpenStudy (anonymous):

same deal as last time, this time it's c/f

OpenStudy (jkbo):

ok than you I understand now

OpenStudy (anonymous):

welcome ^^ .122m yah?